# Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity

### v1

@@ A.3 A note on unbiased estimation and variance reduction
$\left\langle b\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}=b\,\frac{\partial}{\partial Z}\int P(X(T)\mid Z)\,dX=0.$ (A.13)
Appendix B Eligibility traces for STD synapses
@@
B.1 Eligibility as a filtered, parameter-weighted presynaptic spike train
@@
$u_{i}(t)=h_{i}(t,\theta)+\sum_{k=1}^{N}(\epsilon*y_{ik})(t),\qquad(\epsilon*y)(t):=\int_{0}^{t}\epsilon(t-s)\,y(s)\,ds.$ (B.2)
@@
$e_{ij}^{Z}(t)=\frac{\partial u_{i}(t)}{\partial Z_{ij}}=(\epsilon*\psi_{ij}^{Z})(t),\qquad\psi_{ij}^{Z}(t):=\frac{\partial y_{ij}(t)}{\partial Z_{ij}}=\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$ (B.3)
B.2 Exponential EPSP kernel: ODE / event-driven update
For $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the convolution representation (B.3) is equivalent to the linear ODE
$\dot{e}_{ij}^{Z}(t)=-\frac{1}{\tau_{m}}e_{ij}^{Z}(t)+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$ (B.4)
@@ B.2 Exponential EPSP kernel: ODE / event-driven update
$\begin{cases}\text{between spikes:}\quad&e_{ij}^{Z}(t+\Delta t)=e_{ij}^{Z}(t)\,e^{-\Delta t/\tau_{m}},\\ \text{at }t=t_{j}^{f}:\quad&e_{ij}^{Z}(t^{+})=e_{ij}^{Z}(t^{-})+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}.\end{cases}$ (B.5)
B.3 STD dynamics and parameter sensitivities
@@
$s_{ij}^{Z}(t):=\frac{\partial d_{ij}(t)}{\partial Z_{ij}}.$ (B.8)
@@
$s_{ij}^{w_{0}}(t)\equiv 0.$ (B.9)
For $Z_{ij}=U_{ij}$, differentiating (B.7) yields the distributional ODE
$\dot{s}_{ij}^{U}(t)=-\frac{1}{\tau_{d}}s_{ij}^{U}(t)-\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right]x_{j}(t),$ (B.10)
@@
$\begin{cases}\text{between spikes:}\quad&s_{ij}^{U}(t+\Delta t)=s_{ij}^{U}(t)\,e^{-\Delta t/\tau_{d}},\\ \text{at }t=t_{j}^{f}:\quad&s_{ij}^{U}(t^{+})=(1-U_{ij})\,s_{ij}^{U}(t^{-})-d_{ij}(t^{-}).\end{cases}$ (B.11)
@@
$\frac{\partial w_{ij}(t^{-})}{\partial w_{ij}^{0}}$ $=U_{ij}\,d_{ij}(t^{-}),$ (B.12)
$\frac{\partial w_{ij}(t^{-})}{\partial U_{ij}}$ $=w_{ij}^{0}\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right].$ (B.13)
Substituting (B.12) into (B.4) yields an explicit online computation of $e_{ij}^{Z}(t)$ from the simulated presynaptic spikes and STP state variables.
Appendix C Weak-coupling reduction: factorization and closed STD sensitivity dynamics
@@
C.1 Why the score term vanishes at $w=0$
@@
$\begin{split}\left\langle\mathcal{J}[X]\,\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}=\mathcal{J}_{0}\,\left\langle\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}\\ =\mathcal{J}_{0}\,\frac{\partial}{\partial Z_{ij}}\int P(X(T)\mid Z)\,dX=0.\end{split}$ (C.1)
C.2 Compensation formula at $w=0$ and the factorized expectation
@@
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\left\langle\int_{0}^{T}\phi(t)\,\nu_{j}^{0}(t)\,dt\right\rangle_{w=0}.$ (C.2)
@@
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\int_{0}^{T}\nu_{j}^{0}(t)\,\left\langle\phi(t)\right\rangle_{w=0}\,dt.$ (C.3)
@@
$\left\langle e_{ij}^{Z}(t)\right\rangle_{w=0}=\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\nu_{j}^{0}(t^{\prime})\,\frac{\partial}{\partial Z_{ij}}\left\langle w_{ij}(t^{\prime})\right\rangle_{w=0}.$ (C.4)
@@
C.3 Closed dynamics of the mean efficacy and normalized sensitivities
@@
C.3.1 Mean dynamics under Poisson spiking
@@
$\dot{w}(t)=\frac{w^{0}U-w(t)}{\tau_{d}}-Uw(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr),$ (C.6)
@@
$m(t):=\bigl\langle w(t)\bigr\rangle_{X(t)}.$ (C.7)
@@
$\bigl\langle w(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr)\bigr\rangle_{X(t)}=\nu^{0}(t)\,\bigl\langle w(t)\bigr\rangle_{X(t)}=\nu^{0}(t)\,m(t).$ (C.8)
@@
$\dot{m}(t)=\frac{w^{0}U-m(t)}{\tau_{d}}-U\nu^{0}(t)\,m(t).$ (C.9)
C.3.2 Parameter sensitivities and normalized sensitivity functions
Differentiate (C.9) with respect to $w^{0}$ (noting $\nu^{0}(t)$ and $U$ do not depend on $w^{0}$):
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial w^{0}}$ $=\frac{U}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial w^{0}}.$ (C.10)
@@ C.3.2 Parameter sensitivities and normalized sensitivity functions
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial U}$ $=\frac{w^{0}}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial U}-\nu^{0}(t)\,m(t).$ (C.11)
Because (C.9) is linear in $m(t)$ and proportional to $w^{0}$ (given our initial conditions), the solution satisfies $m(t)=w^{0}\,y(t;U)$ and hence
$\frac{\partial m(t)}{\partial w^{0}}=\frac{m(t)}{w^{0}}.$ (C.12)
@@
$f^{w_{0}}(t)$ $:=\frac{1}{U}\,\frac{\partial m(t)}{\partial w^{0}},$ $f^{U}(t)$ $:=\frac{1}{w^{0}}\,\frac{\partial m(t)}{\partial U}.$ (C.13)
Using (C.10) yields
$\dot{f}^{w_{0}}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{w_{0}}(t),$ (C.14)
and using (C.11) together with (C.12) yields
$\dot{f}^{U}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{U}(t)-\nu^{0}(t)\,Uf^{w_{0}}(t).$ (C.15)
@@
Appendix D Analysis of STD sensitivity functions and frequency-domain response
This appendix analyzes the properties of the sensitivity functions $f^{w_{0}}(t)$ and $f^{U}(t)$ and the effective presynaptic learning term $C^{Z}(t)=f^{Z}(t)\nu(t)$, starting from the closed weak-coupling dynamics derived in Appendix C.3 (equivalently, main text Eq. (17)). For notational simplicity, we omit synaptic indices.
D.1 Response to constant input
@@
$f^{w_{0}}_{*}=\frac{1}{1+\tau_{d}\nu U},\quad f^{U}_{*}=\frac{1}{(1+\tau_{d}\nu U)^{2}}.$ (D.1)
@@
$f^{w_{0}}(t)$ $=f^{w_{0}}_{*}+\left[f^{w_{0}}(0)-f^{w_{0}}_{*}\right]e^{-\kappa t}$ (D.2)
$f^{U}(t)$ $=f^{U}_{*}+\bigg[\left(f^{U}(0)-f^{U}_{*}\right)-\nu U\left(f^{w_{0}}(0)-f^{w_{0}}_{*}\right)t\bigg]e^{-\kappa t},$ (D.3)
@@
D.2 Linear response to sinusoidal modulation
@@
$f^{w_{0}}(t)=f^{w_{0}}_{*}+\delta f^{w_{0}}(t),\qquad f^{U}(t)=f^{U}_{*}+\delta f^{U}(t).$ (D.5)
@@
$\dot{\delta f}^{w_{0}}(t)$ $=-\kappa\delta f^{w_{0}}(t)-Uf^{w_{0}}_{*}\delta\nu(t),$ (D.6)
$\dot{\delta f}^{U}(t)$ $=-\kappa\delta f^{U}(t)-\nu_{0}U\delta f^{w_{0}}(t)-U(f^{U}_{*}+f^{w_{0}}_{*})\delta\nu(t),$ (D.7)
@@
$H_{f}^{w_{0}}(\omega)$ $=-\frac{Uf^{w_{0}}_{*}}{\kappa+i\omega},$ (D.8)
$H_{f}^{U}(\omega)$ $=-\frac{U(f^{U}_{*}+f^{w_{0}}_{*})}{\kappa+i\omega}+\frac{\nu_{0}U^{2}f^{w_{0}}_{*}}{(\kappa+i\omega)^{2}}.$ (D.9)
@@
$|H_{f}^{w_{0}}(\omega)|$ $=\frac{Uf^{w_{0}}_{*}}{\sqrt{\kappa^{2}+\omega^{2}}},$ (D.10)
$\phi_{f}^{w_{0}}(\omega)$ $=\pi-\tan^{-1}\left(\frac{\omega}{\kappa}\right).$ (D.11)
@@
D.2.2 Response of the effective presynaptic term.
@@
$\delta C^{Z}(t)\approx f^{Z}_{*}\delta\nu(t)+\nu_{0}\delta f^{Z}(t).$ (D.12)
@@
$H_{C}^{Z}(\omega)=f^{Z}_{*}+\nu_{0}H_{f}^{Z}(\omega).$ (D.13)
D.2.3 Response characteristic for $Z=w_{0}$.
@@
$H_{C}^{w_{0}}(\omega)$ $=f^{w_{0}}_{*}\left(1-\frac{\nu_{0}U}{\kappa+i\omega}\right)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}+i\omega}{\kappa+i\omega}.$ (D.14)
@@
$\phi(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right),$ (D.15)
@@
$\frac{\mathrm{d}\phi}{\mathrm{d}\omega}=\frac{\tau_{d}^{-1}}{\tau_{d}^{-2}+\omega^{2}}-\frac{\kappa}{\kappa^{2}+\omega^{2}}=0.$ (D.16)
@@
$\omega_{*,w_{0}}=\sqrt{\tau_{d}^{-1}\kappa}.$ (D.18)
@@
$H_{C}^{w_{0}}(0)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}}{\kappa}=(\kappa\tau_{d})^{-2}.$ (D.19)
@@
$H_{C}^{w_{0}}(0)=\left.\frac{\mathrm{d}}{\mathrm{d}\nu}\bigl[\nu f_{*}^{w_{0}}(\nu)\bigr]\right|_{\nu=\nu_{0}}=f_{*}^{w_{0}}(\nu_{0})+\nu_{0}\frac{\partial f_{*}^{w_{0}}}{\partial\nu_{0}},$ (D.20)
@@
$\lim_{\omega\to\infty}H_{C}^{w_{0}}(\omega)=\frac{1}{\tau_{d}\kappa}=f^{w_{0}}_{*}.$ (D.21)
@@
$H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}+i\omega}{(\kappa+i\omega)^{2}}\left(-\tau_{d}\kappa^{2}+2\kappa+i\omega\right).$ (D.22)
@@
$r:=\kappa\tau_{d},\qquad x:=\frac{\omega}{\kappa}.$ (D.23)
@@
$H_{C}^{U}(\omega)=\frac{1}{r^{2}}\,\frac{\bigl(\frac{1}{r}+ix\bigr)\bigl((2-r)+ix\bigr)}{(1+ix)^{2}},$ (D.24)
@@
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x).$ (D.25)
@@
$\frac{\mathrm{d}\phi_{U}}{\mathrm{d}x}=\frac{r}{1+r^{2}x^{2}}+\frac{2-r}{(2-r)^{2}+x^{2}}-\frac{2}{1+x^{2}}=0.$ (D.26)
@@
$\omega_{*,U}^{2}=\kappa^{2}\,\frac{-r(r-2)(r-1)+\sqrt{r(r-2)\bigl\{(r-1)^{4}-4\bigr\}}}{r(r+1)}.$ (D.27)
@@
$\lim_{\omega\to 0}H_{C}^{U}(\omega)$ $=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}}{\kappa^{2}}\,\bigl(-\tau_{d}\kappa^{2}+2\kappa\bigr)$ (D.28)
$=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}.$ (D.29)
@@
$f_{*}^{U}(\nu)=\frac{1}{\bigl(1+\tau_{d}\nu U\bigr)^{2}},$ (D.30)
@@
$C_{*}^{U}(\nu)=\nu f_{*}^{U}(\nu)=\frac{\nu}{\bigl(1+\tau_{d}\nu U\bigr)^{2}}.$ (D.31)

### v2

@@ 5.1.3 A note on unbiased estimation and variance reduction
$\left\langle b\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}\\ =b\,\frac{\partial}{\partial Z}\int P(X(T)\mid Z)\,dX=0.$ (S13)
5.2 Eligibility traces for STD synapses
@@
5.2.1 Eligibility as a filtered, parameter-weighted presynaptic spike train
@@
$u_{i}(t)=h_{i}(t,\theta)+\sum_{k=1}^{N}(\epsilon*y_{ik})(t),$ (S15)
$(\epsilon*y)(t)=\int_{0}^{t}\epsilon(t-s)\,y(s)\,ds.$
@@
$e_{ij}^{Z}(t)=\frac{\partial u_{i}(t)}{\partial Z_{ij}}=(\epsilon*\psi_{ij}^{Z})(t),$ (S16)
$\psi_{ij}^{Z}(t)=\frac{\partial y_{ij}(t)}{\partial Z_{ij}}=\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$
5.2.2 Exponential EPSP kernel: ODE / event-driven update
For $\epsilon(t)=e^{-t/\tau_{m}}\Theta(t)$, the convolution representation (S16) is equivalent to the linear ODE
$\dot{e}_{ij}^{Z}(t)=-\frac{1}{\tau_{m}}e_{ij}^{Z}(t)+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}\,x_{j}(t).$ (S17)
@@ 5.2.2 Exponential EPSP kernel: ODE / event-driven update
$\begin{cases}\text{between spikes:}\quad&e_{ij}^{Z}(t+\Delta t)=e_{ij}^{Z}(t)\,e^{-\Delta t/\tau_{m}},\\ \text{at }t=t_{j}^{f}:\quad&e_{ij}^{Z}(t^{+})=e_{ij}^{Z}(t^{-})+\frac{\partial w_{ij}(t^{-})}{\partial Z_{ij}}.\end{cases}$ (S18)
5.2.3 STD dynamics and parameter sensitivities
@@
$s_{ij}^{Z}(t):=\frac{\partial d_{ij}(t)}{\partial Z_{ij}}.$ (S21)
@@
$s_{ij}^{w_{0}}(t)\equiv 0.$ (S22)
For $Z_{ij}=U_{ij}$, differentiating (S20) yields the distributional ODE
$\dot{s}_{ij}^{U}(t)=-\frac{1}{\tau_{d}}s_{ij}^{U}(t)-\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right]x_{j}(t),$ (S23)
@@
$\begin{cases}\text{between spikes:}\quad&s_{ij}^{U}(t+\Delta t)=s_{ij}^{U}(t)\,e^{-\Delta t/\tau_{d}},\\ \text{at }t=t_{j}^{f}:\quad&s_{ij}^{U}(t^{+})=(1-U_{ij})\,s_{ij}^{U}(t^{-})-d_{ij}(t^{-}).\end{cases}$ (S24)
@@
$\frac{\partial w_{ij}(t^{-})}{\partial w_{ij}^{0}}$ $=U_{ij}\,d_{ij}(t^{-}),$ (S25)
$\frac{\partial w_{ij}(t^{-})}{\partial U_{ij}}$ $=w_{ij}^{0}\left[d_{ij}(t^{-})+U_{ij}\,s_{ij}^{U}(t^{-})\right].$ (S26)
Substituting (S25) into (S17) yields an explicit online computation of $e_{ij}^{Z}(t)$ from the simulated presynaptic spikes and STP state variables.
5.3 Weak-coupling reduction: factorization and closed STD sensitivity dynamics
@@
5.3.1 Why the score term vanishes at $w=0$
@@
$\begin{split}\left\langle\mathcal{J}[X]\,\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}\\ =\mathcal{J}_{0}\,\left\langle\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\right\rangle_{w=0}\\ =\mathcal{J}_{0}\,\frac{\partial}{\partial Z_{ij}}\int P(X(T)\mid Z)\,dX=0.\end{split}$ (S27)
5.3.2 Compensation formula at $w=0$ and the factorized expectation
@@
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\left\langle\int_{0}^{T}\phi(t)\,\nu_{j}^{0}(t)\,dt\right\rangle_{w=0}.$ (S28)
@@
$\left\langle\int_{0}^{T}\phi(t)\,dN_{j}(t)\right\rangle_{w=0}=\int_{0}^{T}\nu_{j}^{0}(t)\,\left\langle\phi(t)\right\rangle_{w=0}\,dt.$ (S29)
@@
$\left\langle e_{ij}^{Z}(t)\right\rangle_{w=0}=\int_{0}^{t}dt^{\prime}\,\epsilon(t-t^{\prime})\,\nu_{j}^{0}(t^{\prime})\,\frac{\partial}{\partial Z_{ij}}\left\langle w_{ij}(t^{\prime})\right\rangle_{w=0}.$ (S30)
@@
5.3.3 Closed dynamics of the mean efficacy and normalized sensitivities
@@
Mean dynamics under Poisson spiking
@@
$\dot{w}(t)=\frac{w^{0}U-w(t)}{\tau_{d}}-Uw(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr),$ (S32)
@@
$m(t):=\bigl\langle w(t)\bigr\rangle_{X(t)}.$ (S33)
@@
$\bigl\langle w(t^{-})\delta\bigl(t-t^{\text{spike}}\bigr)\bigr\rangle_{X(t)}=\nu^{0}(t)\,\bigl\langle w(t)\bigr\rangle_{X(t)}=\nu^{0}(t)\,m(t).$ (S34)
@@
$\dot{m}(t)=\frac{w^{0}U-m(t)}{\tau_{d}}-U\nu^{0}(t)\,m(t).$ (S35)
Parameter sensitivities and normalized sensitivity functions
Differentiate (S35) with respect to $w^{0}$ (noting $\nu^{0}(t)$ and $U$ do not depend on $w^{0}$):
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial w^{0}}$ $=\frac{U}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial w^{0}}.$ (S36)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\frac{\partial m(t)}{\partial U}$ $=\frac{w^{0}}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)\frac{\partial m(t)}{\partial U}-\nu^{0}(t)\,m(t).$ (S37)
Because (S35) is linear in $m(t)$ and proportional to $w^{0}$ (given our initial conditions), the solution satisfies $m(t)=w^{0}\,y(t;U)$ and hence
$\frac{\partial m(t)}{\partial w^{0}}=\frac{m(t)}{w^{0}}.$ (S38)
@@
$f^{w_{0}}(t)$ $:=\frac{1}{U}\,\frac{\partial m(t)}{\partial w^{0}},$ $f^{U}(t)$ $:=\frac{1}{w^{0}}\,\frac{\partial m(t)}{\partial U}.$ (S39)
Using (S36) yields
$\dot{f}^{w_{0}}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{w_{0}}(t),$ (S40)
and using (S37) together with (S38) yields
$\dot{f}^{U}(t)$ $=\frac{1}{\tau_{d}}-\left(\frac{1}{\tau_{d}}+U\nu^{0}(t)\right)f^{U}(t)-\nu^{0}(t)\,Uf^{w_{0}}(t).$ (S41)
@@
5.4 Analysis of STD sensitivity functions and frequency-domain response
This appendix analyzes the properties of the sensitivity functions $f^{w_{0}}(t)$ and $f^{U}(t)$ and the effective presynaptic learning term $C^{Z}(t)=f^{Z}(t)\nu(t)$, starting from the closed weak-coupling dynamics derived in Appendix 5.3.3 (equivalently, main text Eq. (17)). For notational simplicity, we omit synaptic indices.
5.4.1 Response to constant input
@@
$f^{w_{0}}_{*}=\frac{1}{1+\tau_{d}\nu U},\quad f^{U}_{*}=\frac{1}{(1+\tau_{d}\nu U)^{2}}.$ (S42)
@@
$f^{w_{0}}(t)$ $=f^{w_{0}}_{*}+\left[f^{w_{0}}(0)-f^{w_{0}}_{*}\right]e^{-\kappa t}$ (S43)
$f^{U}(t)$ $=f^{U}_{*}+\bigg[\left(f^{U}(0)-f^{U}_{*}\right)-\nu U\left(f^{w_{0}}(0)-f^{w_{0}}_{*}\right)t\bigg]e^{-\kappa t},$ (S44)
@@
5.4.2 Linear response to sinusoidal modulation
@@
$f^{w_{0}}(t)=f^{w_{0}}_{*}+\delta f^{w_{0}}(t),\qquad f^{U}(t)=f^{U}_{*}+\delta f^{U}(t).$ (S46)
@@
$\dot{\delta f}^{w_{0}}(t)$ $=-\kappa\delta f^{w_{0}}(t)-Uf^{w_{0}}_{*}\delta\nu(t),$ (S47)
$\dot{\delta f}^{U}(t)$ $=-\kappa\delta f^{U}(t)-\nu_{0}U\delta f^{w_{0}}(t)-U(f^{U}_{*}+f^{w_{0}}_{*})\delta\nu(t),$ (S48)
@@
$H_{f}^{w_{0}}(\omega)$ $=-\frac{Uf^{w_{0}}_{*}}{\kappa+i\omega},$ (S49)
$H_{f}^{U}(\omega)$ $=-\frac{U(f^{U}_{*}+f^{w_{0}}_{*})}{\kappa+i\omega}+\frac{\nu_{0}U^{2}f^{w_{0}}_{*}}{(\kappa+i\omega)^{2}}.$ (S50)
@@
$|H_{f}^{w_{0}}(\omega)|$ $=\frac{Uf^{w_{0}}_{*}}{\sqrt{\kappa^{2}+\omega^{2}}},$ (S51)
$\phi_{f}^{w_{0}}(\omega)$ $=\pi-\tan^{-1}\left(\frac{\omega}{\kappa}\right).$ (S52)
@@
5.4.3 Response of the effective presynaptic term.
@@
$\delta C^{Z}(t)\approx f^{Z}_{*}\delta\nu(t)+\nu_{0}\delta f^{Z}(t).$ (S53)
@@
$H_{C}^{Z}(\omega)=f^{Z}_{*}+\nu_{0}H_{f}^{Z}(\omega).$ (S54)
Response characteristic for $Z=w_{0}$.
@@
$H_{C}^{w_{0}}(\omega)$ $=f^{w_{0}}_{*}\left(1-\frac{\nu_{0}U}{\kappa+i\omega}\right)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}+i\omega}{\kappa+i\omega}.$ (S55)
@@
$\phi(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right),$ (S56)
@@
$\frac{\mathrm{d}\phi}{\mathrm{d}\omega}=\frac{\tau_{d}^{-1}}{\tau_{d}^{-2}+\omega^{2}}-\frac{\kappa}{\kappa^{2}+\omega^{2}}=0.$ (S57)
@@
$\omega_{*,w_{0}}=\sqrt{\tau_{d}^{-1}\kappa}.$ (S59)
@@
$H_{C}^{w_{0}}(0)=\frac{1}{\tau_{d}\kappa}\frac{\tau_{d}^{-1}}{\kappa}=(\kappa\tau_{d})^{-2}.$ (S60)
@@
$H_{C}^{w_{0}}(0)=\left.\frac{\mathrm{d}}{\mathrm{d}\nu}\bigl[\nu f_{*}^{w_{0}}(\nu)\bigr]\right|_{\nu=\nu_{0}}=f_{*}^{w_{0}}(\nu_{0})+\nu_{0}\frac{\partial f_{*}^{w_{0}}}{\partial\nu_{0}},$ (S61)
@@
$\lim_{\omega\to\infty}H_{C}^{w_{0}}(\omega)=\frac{1}{\tau_{d}\kappa}=f^{w_{0}}_{*}.$ (S62)
@@
$H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}+i\omega}{(\kappa+i\omega)^{2}}\left(-\tau_{d}\kappa^{2}+2\kappa+i\omega\right).$ (S63)
@@
$r:=\kappa\tau_{d},\qquad x:=\frac{\omega}{\kappa}.$ (S64)
@@
$H_{C}^{U}(\omega)=\frac{1}{r^{2}}\,\frac{\bigl(\frac{1}{r}+ix\bigr)\bigl((2-r)+ix\bigr)}{(1+ix)^{2}},$ (S65)
@@
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x).$ (S66)
@@
$\frac{\mathrm{d}\phi_{U}}{\mathrm{d}x}=\frac{r}{1+r^{2}x^{2}}+\frac{2-r}{(2-r)^{2}+x^{2}}-\frac{2}{1+x^{2}}=0.$ (S67)
@@
$\omega_{*,U}^{2}=\kappa^{2}\,\frac{-r(r-2)(r-1)+\sqrt{r(r-2)\bigl\{(r-1)^{4}-4\bigr\}}}{r(r+1)}.$ (S68)
@@
$\lim_{\omega\to 0}H_{C}^{U}(\omega)$ $=\frac{1}{(\tau_{d}\kappa)^{2}}\,\frac{\tau_{d}^{-1}}{\kappa^{2}}\,\bigl(-\tau_{d}\kappa^{2}+2\kappa\bigr)$ (S69)
$=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}.$ (S70)
@@
$f_{*}^{U}(\nu)=\frac{1}{\bigl(1+\tau_{d}\nu U\bigr)^{2}},$ (S71)
@@
$C_{*}^{U}(\nu)=\nu f_{*}^{U}(\nu)=\frac{\nu}{\bigl(1+\tau_{d}\nu U\bigr)^{2}}.$ (S72)